OpenStudy (anonymous):

you spin the pointer on the spinner . find the probability that the pointer will stop on a number less then 5 . <3 PLEASE HELP ! <3

OpenStudy (anonymous):

How many total numbers are on the spinner?

OpenStudy (anonymous):

4/6 if there are 6 numbers

OpenStudy (anonymous):

oh yea right its 8 numbers on the spinner .

OpenStudy (anonymous):

Yeah. You need to tell us more about the spinner. How many numbers? Are each of the numbers equally likely, or do some of the numbers have more area?

OpenStudy (anonymous):

So 4/8 = 1/2

OpenStudy (anonymous):

so it would be 5/8?

OpenStudy (anonymous):

then its 4/8

OpenStudy (anonymous):

oh ohkay <3 thanks guys :)

OpenStudy (anonymous):

how many nubers are there

OpenStudy (anonymous):

Assuming equally likely outcomes, the probability is: (number of "desirable" outcomes)/(total number of outcomes) In this case, "desirable" means the outcomes less than 5. Those outcomes are 1, 2 , 3 and 4. We don't count 5 because 5 is not less than 5. The total outcomes are 1, 2, 3, 4, 5, 6, 7, and 8. 4 desirable outcomes / 8 total outcomes = 4/8 = 1/2 = 0.5